Show complete solution. Use equation function.

1)Find the general solution of xdx = 9x^2x -ydy


1
Expert's answer
2021-05-31T18:58:24-0400
"xdx=9x^2dx-ydy"

"\int ydy=\int(9x^2-x)dx"

"\dfrac{1}{2}y^2=3x^3-\dfrac{1}{2}x^2+\dfrac{1}{2}C"

"y^2=6x^3-x^2+C"